OpenStudy (anonymous):

How do i solve y/y+7=2. Pls explain, thanks!

OpenStudy (anonymous):

\[\frac{y}{y+7}=2\iff y=2(y+7)\] is a start

OpenStudy (anonymous):

Thank you!